# On Context Utilization in Summarization with Large Language Models

### v3

@@ 4.2 Can we alleviate the middle curse?
To evaluate loss of performance due to the middle curse in a natural setup, we subsample data points from each of the 5 long-input summarization datasets. We obtain sentences from the (untruncated) source aligned with the reference summary, following the procedure from Section 3.2. Only data points where the start index of the earliest aligned source sentence is at least 1,200 words, are kept, ensuring no salient information at the start. We randomly sample 50 data points from each of Arxiv, PubMed, GovReport and Multi-News, and 25 from SummScreenFD, forming an evaluation dataset of 225 samples which we name MiddleSum. [^2] We will release it alongside our code
@@
$\forall i\in\{1,\ldots,k\},{\bm{y}}_{i}=\text{LLM}({\bm{y}}_{i-1},{\bm{x}}_{i})$ (3)
@@
Results are shown in Fig. 5 (blue and purple bars), with statistics, bigrams analysis and results split over datasets in Appendix E. Both methods show promising results, notably on Mistral-7B for which they improve performance significantly. Across domains, hierarchical and incremental inference are very successful on scientific publications, which we hypothesize is due to the natural division in structured sections of such inputs. Yet, they seem to harm summaries on the other domains.
@@ 4.3 Is scaling context length really useful?
Figure 6: Long-input summarization performance on GovReport with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
Experiments from Section 4.1 confirm that LLMs struggle to summarize information contained in the middle of their context window. This poses issues for long-input summarization: after the initial part with high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens. To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6. We use our longest context LLMs Xgen-7B and Mistral-7B, as well as two LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window. We use two LLMs applying position interpolation: Vicuna-7B-1.5-16k [^3] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^4] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
Results for GovReport in Fig. 6 (and Appendix F for Arxiv and SummScreen) confirm our intuition: all metrics plateau or even decrease (see Mistral-7B) from 4k context window upwards. Two conflicting forces are at play when increasing length: giving more information to the model helps it retrieving key elements to make a summary, while reasoning over a longer context is more challenging. Yet, such a drop for Xgen-7B and Mistral-7B at 8k inference length is concerning. Both position interpolated models show more robustness ; thus these patterns may vary a lot across LLMs. Our results suggest that in the current LLMs inference and evaluation framework, there is no need to exceed 4k tokens in the context window.
@@ 5 Related Work
It is widely acknowledged that LLMs have propelled forward abstractive summarization research (Pu et al., 2023), with their summaries being highly rated by human annotators (Goyal et al., 2022; Zhang et al., 2023b). Liu et al. (2023b) proposes to train smaller models like BART (Lewis et al., 2020) or BRIO (Liu et al., 2022b) with contrastive learning using LLMs like ChatGPT as evaluator providing signal on which generated summary candidate is better. Summary chain-of-thought designs a custom chain-of-thought method which first prompts the LLM to list important facts, then integrates these facts into a coherent summary (Wang et al., 2023c). SummIt utilizes ChatGPT to iteratively write then refine summaries given feedback from an evaluator LLM (Zhang et al., 2023a). Chain-of-density gradually makes GPT-4 generated summaries contain more and more entities while keeping length budget constant, creating more informative albeit a bit less readable summaries (Adams et al., 2023). Ravaut et al. (2022) noticed that data points with higher compression are generally harder to summarize with pre-trained models.
@@
Sun et al. (2021) showed that for Transformer-based models, most recent tokens play a greater role compared to older tokens for next-token prediction. It was later found that for in-context learning, the order of examples within the prompt impacts GPT-3’s performance Liu et al. (2021); Lu et al. (2021). Reliance on positional information affects LLMs capabilities in arithmetic (Shen et al., 2023b), in multiple-choice question-answering (Zheng et al., 2023; Pezeshkpour and Hruschka, 2023), and as text generation evaluators (Wang et al., 2023b) ; making it hard to rank LLMs (Alzahrani et al., 2024). Liu et al. (2023a) was the first to show that LLMs’ performance weakens in the middle of the prompt, yet, how LLMs make use of their full context window remains poorly understood. The passkey retrieval evaluation, which consists in prompting the LLM to recall a complex string or long number inserted in its prompt, is becoming popular recently as a way of verifying LLM’s processing capability at each position Liu et al. (2023a); Jiang et al. (2024). However, this task does not measure position bias on complex, abstract reasoning tasks like summarization. A line of work attempts to solve the middle curse through compressing the prompt Jiang et al. (2023b, c), with very promising results albeit at the cost of prompt fluency. Another approach marginalizes results over different permutations of the input to suppress dependency on input order Tang et al. (2023). Concurrent work to ours also finds that in zero-shot summarization, LLMs tend to prefer lead content Chhabra et al. (2024).
@@ 6 Conclusion
Behind the recent hype around LLMs and their amazing instruction following and content generation capacities, our study showcases a major weakness in abstractive summarization: LLMs suffer from the middle curse and struggle to use information in the middle of their context window. LLMs do not make a consistent use of their context window as they mostly look at the beginning and (to a lesser extent) the end, which at first glance may be hidden by the prevalent lead bias in summarization datasets. Extending context window beyond 4k tokens, which has been an intense area of focus lately, is not justified in the current inference and evaluation setup in abstractive summarization. We benchmarked two alternative inference methods (hierarchical inference and incremental inference) on MiddleSum, an evaluation subset designed to showcase the middle curse. Despite promising results, especially on scientific paper datasets, these methods are far from a silver bullet to the middle curse. We call for a better evaluation of LLMs, which accounts for the salient spans of the source which are effectively being processed.
@@ Appendix A Statistics
Table 5: Performance achieved by the 5 LLMs on each dataset for 5 automatic metrics. # sents is the average number of sentences in generated summaries. Multi-X is short for Multi-XScience, Multi-N is Multi-News dataset, AVG columns represent the average over standard-length and long-input datasets, respectively. SOTA numbers are taken from(Xie et al., 2023) on CNN/DM, from (Zhao et al., 2022b) XSum, Reddit-TIFU and SAMSum, from (Pang et al., 2022) for Arxiv and PubMed, from (Xiong et al., 2022) for GovReport and SummScreenFD, from (Xiao et al., 2022) for Multi-News and Multi-XScience. Due to a lack of reported results for other metrics, we only include ROUGE-2 scores for SOTA models. Best scores (outside of SOTA) are in bold.
@@
Table 6: P-value of a 2-sample Kolmogorov-Smirnov test between the position distribution of bigrams in LLM-generated summaries and bigrams in reference summaries. We round numbers to 3 decimals. Numbers in gray correspond to non-significant differences (p-value above 0.05).
@@ Appendix C Baseline Performance
(a) Llama-2-7B
(b) Llama-2-13B
(c) Xgen-7B
(d) Mistral-7B
Figure 8: Relative location of summary bigrams within the source on MiddleSum for 4 LLMs.
In Table 5, we report zero-shot performance with the 5 LLMs described in Section 2. We note that for standard-length datasets, Flan-UL2 is dominating, while for long-context ones, Llama-2-13B takes the lead.
@@ Appendix D Statistical Significance on RQ1
Appendix E More Results on MiddleSum
@@
In Figs. 7(a), 7(b), 7(c) and 7(d), we repeat the salient bigrams analysis from Fig. 1 for each of the four long-input LLMs on MiddleSum. We note that both hierarchical and incremental inference notably decrease reliance on lead bigrams compared to standard inference.
In Figs. 9, 10, 11, 12 and 13, and in Table 8 (exact numbers), we show reference-based evaluation on each of the 5 subsets of MiddleSum.
Figure 9: Reference-based evaluation on the Arxiv subset of MiddleSum dataset.
Figure 10: Reference-based evaluation on the PubMed subset of MiddleSum dataset.
Figure 11: Reference-based evaluation on the GovReport subset of MiddleSum dataset.
Figure 12: Reference-based evaluation on the SummScreenFD subset of MiddleSum dataset.
Figure 13: Reference-based evaluation on the Multi-News subset of MiddleSum dataset.
Appendix F More Results on Long-Input Tasks
In Figs. 14 and 15, we conduct the analysis on increasingly longer input context from Fig. 6 on Arxiv and SummScreen, respectively.
Figure 14: Long-input summarization performance on Arxiv with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
Figure 15: Long-input summarization performance on SummScreen with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.

### v4

@@ 4.2 Can we alleviate the middle curse?
To evaluate the loss of performance due to the middle curse in a natural setup, we subsample data points from each of the 5 long-input summarization datasets. We obtain sentences from the (untruncated) source aligned with the reference summary, following the procedure from Section 3.2. Only data points where the start index of the earliest aligned source sentence is at least 1,200 words, are kept, ensuring no salient information at the start. We randomly sample 50 data points from each of Arxiv, PubMed, GovReport and Multi-News, and 25 from SummScreenFD, forming an evaluation dataset of 225 samples which we name MiddleSum.
@@
${\bm{y}}_{i}$ $=\text{LLM}({\bm{y}}_{i-1},{\bm{x}}_{i})\forall i\in\{1,\ldots,k\}$ (3)
${\bm{y}}$ $={\bm{y}}_{k}$ (4)
Figure 6: Long-input summarization performance on Arxiv (top) and GovReport (bottom) with 5 LLMs and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7B and Mistral-7B cannot infer beyond 8k tokens.
@@
Results are shown in Fig. 5 (blue and purple bars), with detailed numbers in Appendix F. We also compare to a baseline consisting in adding the prompt Please also pay attention to the middle section of the input when constructing the summary, which we refer to as the Focus prompt (brown bars). Both alternative methods show promising results on open-source LLMs, notably on Mistral-7B for which they improve performance significantly. However, they are not successful and lag behind Focus Prompt with GPT-3.5. Across domains (see Table 8), hierarchical and incremental inference are very effective on scientific publications, which we hypothesize is due to the natural division in structured sections of such inputs. Yet, they seem to harm summaries on the other domains.
@@ 4.3 Is scaling context length really useful?
Experiments from Section 4.1 confirm that LLMs struggle to summarize information contained in the middle of their context window. This poses issues for long-input summarization: after the initial part with (usually) high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens. To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.444Our hardware does not allow us to exceed 12k tokens. We use our longest context LLMs {Xgen-7B, Mistral-7B, GPT-3.5} ; as well as two open-source LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window. We use Vicuna-7B-1.5-16k [^5] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^6] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
Results on Arxiv and GovReport in Fig. 6 confirm our intuition: all metrics plateau or even decrease (see Mistral-7B) from 4k context window upwards. Two conflicting forces are at play when increasing length: giving more information to the model helps it retrieve key elements further to make a richer summary, while at the same time reasoning over a longer context is more challenging. Yet, such a drop for Xgen-7B and Mistral-7B at 8k inference length is concerning. Both position interpolated models show more robustness ; while GPT-3.5 seems to plateau at 8k tokens. Our results suggest that in the current LLMs inference and evaluation framework, there is no need to exceed 4k tokens in the context window for open-source model.
4.4 Does the decoding method impact the middle curse?
We now turn our attention to the process controlling summary generation. While we had sampled all summaries with top-k sampling with $k=50$ and temperature $T=0.3$ so far ; we now also experiment with greedy decoding, and top-p sampling (where we use $p=0.95$ and temperature $T=1.0$).
Figure 7: Summary bigrams (top) and aligned source sentences (bottom) distribution on Arxiv and GovReport for Llama-2-7B and XGen-7B, for several decoding strategies.
In Fig. 7, we reproduce the salient bigrams and sentences alignment experiments from Section 3.1 and Section 3.2, respectively, with the aforementioned decoding methods. As seen, the decoding method does not affect position bias: for all setups, the LLMs show similar patterns as with our previous default decoding method. We conclude that the middle curse is independent from the decoding method.
@@ 5 Related Work
It is widely acknowledged that LLMs have propelled forward abstractive summarization research (Pu et al., 2023), with their summaries being highly rated by human annotators (Goyal et al., 2022; Zhang et al., 2023b). Liu et al. (2023d) proposes to train smaller models like BART (Lewis et al., 2020) or BRIO (Liu et al., 2022b) with contrastive learning using LLMs like ChatGPT as evaluator providing signal on which generated summary candidate is better. Summary chain-of-thought designs a custom chain-of-thought method which first prompts the LLM to list important facts, then integrates these facts into a coherent summary (Wang et al., 2023c). SummIt utilizes ChatGPT to iteratively write then refine summaries given feedback from an evaluator LLM (Zhang et al., 2023a). Chain-of-density gradually makes GPT-4 generated summaries contain more and more entities while keeping length budget constant, creating more informative albeit a bit less readable summaries (Adams et al., 2023). Ravaut et al. (2022) noticed that data points with higher compression are generally harder to summarize with pre-trained models.
@@
Sun et al. (2021) showed that for Transformer-based models, most recent tokens play a greater role compared to older tokens for next-token prediction. It was later found that for in-context learning, the order of examples within the prompt impacts GPT-3’s performance Liu et al. (2022a); Lu et al. (2021). Reliance on positional information affects LLMs capabilities in arithmetic (Shen et al., 2023b), in multiple-choice question-answering (Zheng et al., 2023; Pezeshkpour and Hruschka, 2023), and as text generation evaluators (Wang et al., 2023b) ; making it hard to rank LLMs (Alzahrani et al., 2024). Liu et al. (2023a) were the first to show that LLMs’ performance weakens in the middle of the prompt (the middle curse), yet, how LLMs make use of their full context window remains poorly understood. The passkey retrieval evaluation, which consists in prompting the LLM to recall a complex string or long number inserted in its prompt, is becoming popular recently as a way of verifying LLM’s processing capability at each position Liu et al. (2023a); Jiang et al. (2024). However, this task does not measure position bias on complex, abstract reasoning tasks like summarization. A line of work attempts to solve the middle curse through compressing the prompt Jiang et al. (2023b, c), with very promising results albeit at the cost of prompt fluency. Another approach marginalizes results over different permutations of the input to suppress dependency on input order Tang et al. (2023). Concurrent work to ours also finds that in zero-shot summarization, LLMs tend to prefer lead content Chhabra et al. (2024).
@@ 6 Conclusion
Behind the recent hype around LLMs and their amazing instruction following and content generation capabilities, our study showcases a major weakness in abstractive summarization: LLMs suffer from the middle curse and struggle to use information in the middle of their context window. LLMs do not make a consistent use of their context window as they mostly look at the beginning and (to a lesser extent) the end, which at first glance may be hidden by the prevalent lead bias in summarization datasets. Extending context window beyond 4k tokens, which has been an intense area of focus lately, is not justified in the current inference and evaluation setup in abstractive summarization. We benchmarked two alternative inference methods (hierarchical inference and incremental inference) on MiddleSum, an evaluation subset designed to showcase the middle curse. Despite encouraging results on scientific paper datasets, these methods are far from a silver bullet to the middle curse. We call for a better evaluation of LLMs, which accounts for the salient spans of the source which are effectively being processed.
@@
Acknowledgements
This research was supported by the SINGA scholarship and partially supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.
We thank anonymous ARR Reviewers 1 and 3 for their great review, which have led us to include plenty more important results enriching our analysis. We also thank the AC for their constructive meta-review. We thank Florian Le Bronnec and Ruochen Zhao for their helpful proof-reading of the draft.
@@ Appendix A Statistics
[…]
ROUGE-2 13.17 8.19 5.17 14.47 5.37 9.27 13.78 13.90 17.94 6.55 12.28 12.89
BERTScore 87.26 87.90 86.38 89.79 83.82 87.03 84.14 84.29 85.99 86.13 86.02 85.31
A3CU 26.72 15.21 13.37 35.85 7.90 19.81 17.98 18.76 19.32 12.69 22.95 18.34
SummaC 35.14 23.22 26.00 24.92 37.56 29.37 50.11 47.36 73.23 40.58 48.91 52.04
GPT-3.5 4.12 4.58 3.08 3.81 4.61 4.04 4.41 4.36 4.03 3.39 4.14 4.07
Table 5: Performance achieved by the LLMs on each dataset for all 5 metrics. # sents is the average number of sentences in generated summaries. Multi-X is short for Multi-XScience, Multi-N is Multi-News dataset, AVG columns represent the average over standard-length and long-input datasets, respectively. SOTA numbers are taken from (Xie et al., 2023) on CNN/DM, from (Zhao et al., 2022b) XSum, Reddit-TIFU and SAMSum, from (Pang et al., 2022) for Arxiv and PubMed, from (Xiong et al., 2023) for GovReport and SummScreenFD, from (Xiao et al., 2022) for Multi-News and Multi-XScience. Due to a lack of reported results for other metrics, we only include ROUGE-2 scores for SOTA models. Best scores (outside of SOTA) are in bold.
Appendix B GPT-3.5 Evaluation
@@
GPT-3.5 0.000 0.000 0.000 0.135 0.000 0.000 0.000 0.000 0.000 0.000
Table 6: P-value of a 2-sample Kolmogorov-Smirnov test between the position distribution of bigrams in LLM-generated summaries and bigrams in reference summaries. We round numbers to 3 decimals. Numbers in gray correspond to non-significant differences (p-value above 0.01).
@@ Appendix C Baseline Performance
Figure 9: Relative location of summary bigrams within the source on MiddleSum for Llama-2-13B with different inference methods.
In Table 5, we report zero-shot performance with the 6 LLMs described in Section 2. We note that for standard-length datasets, Flan-UL2 is dominating, perhaps due to its better instruction-tuning ; while for long-context ones, GPT-3.5 takes the lead. GPT-3.5 is always consistently ahead on the GPT-3.5 metric, showing the preference of the LLM for its own generation when used as an evaluator.
@@ Appendix D Statistical Significance on RQ1
Appendix E Results with base Models
Figure 10: Summary bigrams (top) and aligned source sentences (bottom) distribution for Llama-2-7B and Llama-2-13B on the long-input datasets, both with base (dashed lines) and chat (full lines) models.
In Fig. 10, we repeat the analysis from Section 3.1 (bigrams alignment) and Section 3.2 (sentences alignment) but this time using the base model versions (not instruction-tuned) of Llama-2-7B and Llama-2-13B. As we can see, the position bias is even stronger for base models than for chatbot models, suggesting that this bias is acquired during pre-training.
Appendix F More Results on MiddleSum
@@
In Fig. 9, we repeat the salient bigrams analysis from Fig. 1 for Llama-2-13B on MiddleSum. We note that both hierarchical and incremental inference notably decrease reliance on lead bigrams compared to standard inference, while the simple Focus Prompt does not.
In Table 8, we show reference-based evaluation on each of the 5 subsets of MiddleSum.
@@
GPT-3.5 ROUGE-2 Standard 12.33 12.99 12.78 16.16 6.05 10.53
Focus prompt 12.17 12.61 12.56 15.95 6.35 10.46
Hierarchical 8.59 9.58 10.27 10.30 4.33 6.36
Incremental 12.04 11.85 12.17 18.46 5.38 9.03
[…]
Focus prompt 84.80 83.76 83.81 85.50 86.28 85.39
Hierarchical 84.27 83.34 83.62 85.30 85.37 84.28
Incremental 84.43 83.48 83.74 85.41 85.76 84.44
[…]
Focus prompt 15.07 14.51 15.59 12.34 13.63 18.54
Hierarchical 12.24 12.54 13.01 13.29 9.03 11.74
Incremental 12.08 13.59 13.24 10.90 9.05 12.08